티15-) Certam amount ofpure hydrogen chloride psisdissolved=water and-ted with-AgNO witine, lfthe mass of precipitate formed from this reaction is 9.730いitthe volnt ofFb-chimte pas under standard temperature and pressure? Assume ideal gas behavice (Gasont 0m mt Atomic masses Ag 107.9 gimol, CI-3545mol MARkS ang 9 (10 points) There is an unknown compound containing m oxygen ga, 30.20 g ofCO, (44 01 g/mol) and 14.30 g of HO(i802 Emol) are prodacod from this proces is the empirical formula of the compound? C, H, and O) When 12.10 g of this compound and Insti tute
Answer #1

9) .mass of sample = 12.10 grams

mass of CO2 = 30.20 grams

molar mass of CO2 = 44 gram/mole

% by mass of C= 12/44 xmass of CO2/mass of sample x 100

                              = 12/44x 30.20/12.10 x100 = 68.069%

% by mass of H = 2/18 x mass of H2O/mass of sample x 100

                              = 2/18 x 14.80/12.10 x 100= 13.59%

% by mass of H= 13.59%

% by mass of O= 100 -[68.069 +13.59]

% by mass of O = 18.341%

Element                     % by mass            atomic wt          Relative number                          simple ratio

C                                  68.069                        12         68.069/12 = 5.672                     5.672/1.146= 4.94 = 5.0..

H                                 13.59                      1.0             13.59/1.0 = 13.59                   13.59/1.146 = 11.85= 12

O                                18.341                     16              18.341/16 = 1.146                    1.146/1.146 = 1.0

Emperical formula = C5H12O.